Shift report 20020115 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Jan 15 2002 - 08:01:26 EST

  • Next message: Jens Jørgen Gaardhøje: "brahms web agenda"

    ------------------------------------------------------------
    Date: Tue Jan 15 08:01:26 2002, Shift: 00:00-08:00
    Supervisor: Mads, with: 
    Log book#12 pages 62 - 62
    
    Shift Summary:
    ------------------------------------------------------------
    There has been trouble with the DAQ throughout the night. And data has only been taken from the MRS. As was done during the end of the previous shift. 
    
    There was a store that lasted until 230, at which datarun 6504 was taken. (mrs only)
    
    The beam got up again at 0530, at wich time i experimented with taking data from the FFS. Evererything looked ok to me at that time, but the beam was suddenly lost at 0540, due to  'permit link interlock'. 
    
    Pedestal runs 6405, 6406,6408 were taken 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Jan 15 2002 - 08:02:29 EST